OpenStudy (anonymous):

Find three solutions of the equation. y = –2x

OpenStudy (anonymous):

Simple plug in 3 values for x Pick: 0, 1, 2

OpenStudy (anonymous):

Once you have done that plug and find out what Y is

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

replace x with 0 and write it down.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

I'm telling you to replace x with zero. Write down what you already wrote but instead of X put a zero.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

Can you write it down?

OpenStudy (anonymous):

\[x=0\]

OpenStudy (anonymous):



OpenStudy (anonymous):

y = –2*x y = –2*0 What's 0 times -2?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

umm no try again

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Yes so when x=0 the solution is 0 now try with x=1

OpenStudy (anonymous):



OpenStudy (anonymous):

ok when x= 1 for y = –2*x you get y = –2*1 What is -2 times 1

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Yes so when x= 1 the solution is -2 Now try x=2 ( This is the last one!!!)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

idk that 1

OpenStudy (anonymous):

X= 2 y = –2*x y = –2*2 What is -2 times 2?

OpenStudy (anonymous):

(0, 0), (3, –8), (–2, –4) (1, –2), (0, 0), (–1, 1) (–2, 4), (1, –2), (0, 0) (–2, 4), (1, –2), (2, –3)

OpenStudy (anonymous):

You have done two already and from the looks of it seems like the answer should be the second one. Good Job!!